# publications.bib

### v3

@@ 1. Introduction
$h_{*}^{\rightarrow}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})\simeq R\Gamma_{c}(\mathcal{S}_{K^{p},C},\overline{\mathbb{F}}_{\ell})\simeq\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathcal{S}_{K^{p}K_{p},C},\overline{\mathbb{F}}_{\ell})\simeq\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K^{p}K_{p},C})$
@@
has a filtration with graded pieces isomorphic to $h^{\rightarrow}_{*}h^{\leftarrow*}(j_{b!}(V_{b}))$ for varying $b\in B(G,\mu)$, where $h^{\rightarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{G(\mathbb{Q}_{p})}]$ is the structure map quotiented by $G(\mathbb{Q}_{p})$. Here the second isomorphism follows since taking compactly supported cohomology respects taking limits of spaces, and the third isomorphism is a standard comparison result due to Huber [Hub, Theorem 3.5.13].
@@
Now, via the Bialynicki-Birula isomorphism, the flag variety $[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$ identifies with an open substack of $\mathrm{Hck}_{G,\leq\mu}$ for the fixed minuscule $\mu$. In particular, under this relationship the maps $h^{\rightarrow}_{\mu}$ and $h^{\leftarrow}_{\mu}$ identify with $h^{\rightarrow}$ and $h^{\leftarrow}$, and therefore we can relate the graded pieces of the excision filtration to Hecke operators. We write
@@
for the complex of $G(\mathbb{Q}_{p})\times J_{b}(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-modules defined by the compactly supported cohomology of this tower. Here $\overline{\mathbb{F}}_{\ell}(d_{b})$ is the sheaf with $J_{b}(\mathbb{Q}_{p})$-action defined as in [Ko, Lemma 7.4].
@@
The complex $R\Gamma_{c}(\mathcal{S}_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ has a filtration as a complex of $G(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-representations with graded pieces isomorphic to $j_{1}^{*}(T_{\mu}j_{b!}(V_{b}))[-d](-\frac{d}{2})$. More specifically, the graded pieces are isomorphic to
@@
as $G(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-modules.
@@
When the Shimura variety is compact, we have that $R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})\simeq R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$, and this recovers precisely [Ko, Theorem 7.1].
We now apply our localization functor $(-)_{\phi_{\mathfrak{m}}}:\mathrm{D}(\mathrm{Bun}_{G})\rightarrow\mathrm{D}(\mathrm{Bun}_{G})_{\phi_{\mathfrak{m}}}$ for a generic maximal ideal $\mathfrak{m}$ to get a complex $R\Gamma_{c}(\mathcal{S}_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\in\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$, which we view as a sheaf on $\mathrm{Bun}_{G}$ by !!! extending along the neutral strata. After applying $R\Gamma(K_{p}^{\mathrm{hs}},-)$, this agrees with $R\Gamma_{c}(\mathcal{S}_{K^{p}K_{p}^{\mathrm{hs}},C},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$, the usual localization under the unramified Hecke algebra, which is the object we want to study. This in turn admits a filtration by $R\Gamma(K_{p}^{\mathrm{hs}},(j_{1}^{*}T_{\mu}j_{b!}(V_{b})))_{\phi_{\mathfrak{m}}}[-d](\frac{-d}{2})$. However, now we know, by the direct sum decomposition of $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$ described above, that the natural map $j_{b!}(V_{b})\rightarrow j_{b*}(V_{b})$ is an isomorphism after applying $(-)_{\phi_{\mathfrak{m}}}$. Moreover, one only has interesting contributions coming from the unramified elements $B(G,\mu)_{\mathrm{un}}$. In particular, we can deduce the following Corollary.
@@
Suppose $(\mathbf{G},X)$ is a PEL datum of type $A$ or $C$ such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of simple groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions, the complex $R\Gamma_{c}(\mathcal{S}_{K,C},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}\simeq R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K,C},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ breaks up as a direct sum
$\bigoplus_{b\in B(G,\mu)_{\mathrm{un}}}(R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\mathfrak{m}}\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell}))[2d_{b}]$
of $H_{K_{p}^{\mathrm{hs}}}\times W_{E_{\mathfrak{p}}}$-modules.
@@
As we will explain more in §6.1, in the case that the unique basic element $b\in B(G,\mu)_{\mathrm{un}}$ is unramified, the contribution of the corresponding summand to middle degree cohomology serves as a generic fiber analogue of the description of the middle degree cohomology on the special fiber of the integral model at hyperspecial level, as provided in [XZ, Theorem 1.1.4].
As a consequence, we deduce our main Theorem, by combining Theorem 1.7 with the fact that $R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})\in\mathrm{D}^{\leq d_{b}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, by Artin vanishing.
Theorem 1.16.
Suppose $(\mathbf{G},X)$ is a PEL datum of type $A$ or $C$ such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of simple groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions then Conjecture 1.2 is true.
Remark 1.17.
This notably allows one to relax the assumption in [CS1, CS2, Ko, San] that the prime $p$ splits in $F$, answering a question of Caraiani.
We can also easily deduce the result for some abelian type Shimura varieties, such as Hilbert modular varieties, from the above result, which recovers work of Caraiani-Tamiozzo [CT] (See Corollary 5.3).
Corollary 1.18.
Suppose $(\mathbf{G},X)$ is an abelian-type Shimura datum which has an associated PEL-type datum $(\mathbf{G}_{1},X_{1})$ of type $A$ or $C$ such that $\mathbf{G}_{1,\mathbb{Q}_{p}}$ is a product of simple groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions. Then Conjecture 1.2 is true.
@@ Acknowledgements
We would like to thank Ana Caraiani, Jean-François Dat, David Hansen, Naoki Imai, Teruhisa Koshikawa, and Chris Skinner for helpful discussions pertaining to this work. Special thanks go to Mafalda Santos for sharing with us the results of her thesis, Peter Scholze for encouraging us to avoid working with the good reduction locus by directly describing the fibers of the Hodge-Tate period morphism, Matteo Tamiozzo for comments and corrections on an earlier draft, as well as suggestions for the arguments in §5.2, and Mingjia Zhang for very helpful discussions and filling in several gaps in the arguments used in §3, as well as several comments and corrections. This project was carried out while the second author was at the Max Planck Institute for Mathematics in Bonn and she thanks them for their hospitality and financial support.
@@
  • We write $\mathbb{Q}_{p}$ for the $p$-adic numbers, and ˘$\breve{\mathbb{Q}}_{p}$ for the completion of the maximal unramified extension with Frobenius $\sigma$.
  • We let $\overline{\mathbb{F}}_{\ell}$ denote the algebraic closure of the finite field $\mathbb{F}_{\ell}$. We fix a choice of square root of $p$ in $\overline{\mathbb{F}}_{\ell}$ and define all half Tate twists and square roots of the norm character with respect to this choice.
  • For $L/\mathbb{Q}_{p}$ a finite extension, we write $\breve{L}:=L\breve{\mathbb{Q}}_{p}$ for the compositum of $L$ with the maximal unramified extension and $W_{L}$ for the Weil group of $L$. We let $\mathrm{WD}_{L}:=W_{L}\times\mathrm{SL}(2,\overline{\mathbb{Q}}_{\ell})$ denote the Weil-Deligne group of $L$.
@@
  • We fix an isomorphism $j:\overline{\mathbb{Q}}_{p}\xrightarrow{\simeq}\mathbb{C}$. Consider the induced embedding ¯¯$\overline{\mathbb{Q}}\rightarrow\overline{\mathbb{Q}}_{p}$ this gives a finite place $\mathfrak{p}$ of $E$. We write $E_{\mathfrak{p}}$ for the completion at $\mathfrak{p}$.
@@
  • If $G$ is unramified then we let $K_{p}^{\mathrm{hs}}\subset G(\mathbb{Q}_{p})$ be a choice of hyperspecial subgroup. We set $H_{K_{p}^{\mathrm{hs}}}:=\overline{\mathbb{F}}_{\ell}[K_{p}^{\mathrm{hs}}\backslash G(\mathbb{Q}_{p})/K_{p}^{\mathrm{hs}}]$ to be the unramified Hecke algebra with $\overline{\mathbb{F}}_{\ell}$-coefficients.
@@
  • For $b\in B(G)$, we write $J_{b}$ for the $\sigma$-centralizer of $b$.
  • For $\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$, we let $B(G,\mu)$ be the $\mu$-admissible locus (as defined in [RV, Definition 2.3]).
@@
  • For $S\in\mathrm{Perf}$, let $X_{S}$ denote the relative schematic Fargues-Fontaine curve over $S$.
@@
  • For $b\in B(G)$, we write $\mathcal{E}_{b}$ for the associated $G$-bundle on $X$.
@@
  • To a diamond or $v$-stack $X$ over $\ast$, we write $\mathrm{D}(X,\overline{\mathbb{F}}_{\ell})$ for the category of étale $\overline{\mathbb{F}}_{\ell}$-sheaves, as defined in [Ecod]. We let $\mathrm{D}^{\mathrm{ULA}}(X,\overline{\mathbb{F}}_{\ell})$ denote the full subcategory of ULA sheaves over $\ast$.
  • For an Artin $v$-stack $X$ and $\Lambda\in\{\overline{\mathbb{F}}_{\ell},\overline{\mathbb{Z}}_{\ell},\overline{\mathbb{Q}}_{\ell}\}$, we write $\mathrm{D}_{\blacksquare}(X,\Lambda)$ for the condensed $\infty$-category of solid $\overline{\mathbb{F}}_{\ell}$-sheaves on $X$, and write $\mathrm{D}_{\mathrm{lis}}(X,\Lambda)\subset\mathrm{D}_{\blacksquare}(X,\Lambda)$ for the full sub-category of $\Lambda$-lisse-étale sheaves, as defined in [FS, Chapter VII].
  • If $X$ is an Artin $v$-stack ([FS, Definition IV.V.1]) admitting a separated cohomologically smooth surjection $U\rightarrow X$ from a locally spatial diamond $U$ such that the étale site has a basis with bounded $\ell$-cohomological dimension (which will always be the case for our applications) then we will regard $\mathrm{D}(X,\overline{\mathbb{F}}_{\ell})$ as a condensed $\infty$-category via the identification $\mathrm{D}_{\mathrm{lis}}(X,\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(X,\overline{\mathbb{F}}_{\ell})$ when viewed as objects in $\mathrm{D}_{\blacksquare}(X,\overline{\mathbb{F}}_{\ell})$ [FS, Proposition VII.6.6].
@@
  • If $E$ denotes the splitting field of $G$ then the action of $W_{\mathbb{Q}_{p}}$on $\hat{G}$ factors through $Q:=W_{\mathbb{Q}_{p}}/W_{E}$. We let $\phantom{}{}^{L}G:=\hat{G}\rtimes Q$ denote the $L$-group.
@@
  • For $\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$, we write $V_{\mu}\in\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\hat{G})$ (resp. $\mathcal{T}_{\mu}\in\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\hat{G})$) for the usual highest weight representation (resp. highest weight tilting module, as in [Don]) of highest weight $\mu$.
  • To any condensed $\infty$-category $\mathcal{C}$, we write $\mathcal{C}^{BW^{I}_{\mathbb{Q}_{p}}}$ for the category of objects with continuous $W_{\mathbb{Q}_{p}}^{I}$-action, as defined in [FS, Section IX.1].
  • For any separated $v$-stack, $X\rightarrow\mathop{\rm Spa}(K,\mathcal{O}_{K})$ where $\mathop{\rm Spa}(K,\mathcal{O}_{K})$ is a non-archimedean field, we write $\overline{X}$ for the canonical compactification of $X$ with respect to the structure map ([Ecod, Proposition 18.6], [Hub, Theorem 5.15]).
  • For a reductive group $H/\mathbb{Q}_{p}$, we write $\mathrm{D}(H(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ for the unbounded derived category of smooth $\overline{\mathbb{F}}_{\ell}$-representations.
  • For an analytic adic space $X$, we will often abuse notation and use $X$ to also denote the diamond $X^{\diamond}$ attached to it (as defined in [SW, Lecture X]).

### v5

@@ 1.0.2. Proof Sketch of the Main Theorem
$h_{*}^{\rightarrow}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})\simeq R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ $\simeq\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p}K_{p},C},\overline{\mathbb{F}}_{\ell})$
$\simeq\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K^{p}K_{p},C},\overline{\mathbb{F}}_{\ell})$
@@
has a filtration with graded pieces isomorphic to $Rh^{\rightarrow}_{*}h^{\leftarrow*}j_{b!}(V_{b})$ for varying $b\in B(G,\mu)$, where $h^{\rightarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{G(\mathbb{Q}_{p})}]$ is the structure map quotiented by $G(\mathbb{Q}_{p})$, which we note is proper. Here the second isomorphism follows since taking compactly supported cohomology respects taking limits of spaces, and the third isomorphism is a standard comparison result due to Huber [32, Theorem 3.5.13].
@@
Now, via the Bialynicki-Birula isomorphism, the flag variety $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ identifies with an open substack of $\mathrm{Hck}_{G,\leq\mu}$ for the fixed minuscule $\mu$. In particular, under this relationship the maps $h^{\rightarrow}_{\mu}$ and $h^{\leftarrow}_{\mu}$ identify with $h^{\rightarrow}$ and $h^{\leftarrow}$, and therefore we can relate the graded pieces of the excision filtration to Hecke operators. We write
@@
for the complex of $G(\mathbb{Q}_{p})\times J_{b}(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-modules defined by the compactly supported cohomology of this tower. Here $\overline{\mathbb{F}}_{\ell}(d_{b})$ is the sheaf with $J_{b}(\mathbb{Q}_{p})$-action defined as in [37, Lemma 7.4].
@@
Suppose that $(\mathbf{G},X)$ is a Shimura datum of PEL type AC satisfying Assumption 1.13 then the complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ has a filtration as a complex of $G(\mathbb{Q}_{p})$-representations with graded pieces isomorphic to $j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d]$. More specifically, the graded pieces are isomorphic to
@@
as $G(\mathbb{Q}_{p})$-modules.
@@
When the Shimura variety is compact, so that $\mathrm{Ig}^{b,*}=\mathrm{Ig}^{b}$ and correspondingly that $R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})\simeq R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$, and this recovers precisely [37, Theorem 7.1].
We now apply our localization functor $(-)_{\phi_{\mathfrak{m}}}:\mathrm{D}(\mathrm{Bun}_{G})\rightarrow\mathrm{D}(\mathrm{Bun}_{G})_{\phi_{\mathfrak{m}}}$ for a generic maximal ideal $\mathfrak{m}$ to get a complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\in\mathrm{D}^{\mathrm{adm}}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$, which we view as a sheaf on $\mathrm{Bun}_{G}$ by !! extending along the neutral strata. Now we know, by the direct sum decomposition of $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$ described in (2), that the natural map $j_{b!}(V_{b})\rightarrow Rj_{b*}(V_{b})$ is an isomorphism after applying $(-)_{\phi_{\mathfrak{m}}}$. Moreover, one only has interesting contributions coming from the unramified elements $B(G,\mu)_{\mathrm{un}}$. In particular, we can deduce the following Corollary of Theorem 1.15.
@@
(Theorem 5.1) Suppose $(\mathbf{G},X)$ is a PEL datum of type AC satisfying Assumption 1.13 such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions and let $\phi_{\mathfrak{m}}$ be a generic parameter of the form described in Theorem 1.11. Then the complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$ breaks up as a direct sum
$\bigoplus_{b\in B(G,\mu)_{\mathrm{un}}}(R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\phi_{\mathfrak{m}}}\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell}))[2d_{b}]$
of $G(\mathbb{Q}_{p})$-modules.
@@
As we will explain more in §6.1, in the case that the unique basic element $b\in B(G,\mu)_{\mathrm{un}}$ is unramified, the contribution of the corresponding summand to middle degree cohomology serves as a generic fiber analogue of the description of the middle degree cohomology on the special fiber of the integral model at hyperspecial level, as provided in [67, Theorem 1.1.4].
It is now easy to see that we obtain from this Theorem 1.8, by combining Theorem 1.11 with the fact that $R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})\in\mathrm{D}^{\leq d_{b}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$, by Artin vanishing applied to the inverse limit defining the partial minimal compactification $\mathrm{Ig}^{b,*}$ (See Proposition 3.6 for details) using a Hochschild-Serre style argument.
@@ Acknowledgements
We would like to thank Ana Caraiani, Jean-François Dat, David Hansen, Naoki Imai, Teruhisa Koshikawa, and Chris Skinner for helpful discussions pertaining to this work. Special thanks go to Mafalda Santos for sharing with us the results of her thesis, Peter Scholze for encouraging us to avoid working with the good reduction locus by directly describing the fibers of the Hodge-Tate period morphism, Matteo Tamiozzo for comments and corrections on an earlier draft, as well as suggestions for the arguments in §5.2, and Mingjia Zhang for very helpful discussions and filling in several gaps in the arguments used in §3, as well as several comments and corrections. We would also like the referees for carefully reading the paper and supplying several useful comments and corrections. This project was carried out while the second author was at the Max Planck Institute for Mathematics in Bonn and she thanks them for their hospitality and financial support.
@@
  • We write $\mathbb{Q}_{p}$ for the $p$-adic numbers, and $\breve{\mathbb{Q}}_{p}$ for the completion of the maximal unramified extension with Frobenius $\sigma$.
  • We let $\overline{\mathbb{F}}_{\ell}$ denote the algebraic closure of the finite field $\mathbb{F}_{\ell}$. We fix a choice of square root of $p$ in $\overline{\mathbb{F}}_{\ell}$ and normalize all half Tate twists and square roots of the norm character with respect to this choice. In particular, this how we normalize the geometric Satake correspondence of [16, Chapter VI] and in turn the Fargues-Scholze correspondence.
  • For $L/\mathbb{Q}_{p}$ a finite extension, we write $\breve{L}:=L\breve{\mathbb{Q}}_{p}$ for the compositum of $L$ with the maximal unramified extension and $W_{L}$ for the Weil group of $L$.
@@
  • We fix an isomorphism $j:\overline{\mathbb{Q}}_{p}\xrightarrow{\simeq}\mathbb{C}$. Consider the induced embedding $\overline{\mathbb{Q}}\rightarrow\overline{\mathbb{Q}}_{p}$ this gives a finite place $\mathfrak{p}$ of $E$. We write $E_{\mathfrak{p}}$ for the completion at $\mathfrak{p}$.
@@
[…]
@@
  • For $b\in B(G)$, we write $J_{b}$ for the $\sigma$-centralizer of $b$. We note that this depends on a choice of representative in the $\sigma$-conjugacy class attached to $b$, but the isomorphism class does not, as a result we will suppress this technicality.
  • For $\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$, we let $B(G,\mu)$ be the $\mu$-admissible locus (as defined in [52, Definition 2.3]).
@@
  • For $S\in\mathrm{Perf}$, let $X_{S}$ denote the relative schematic Fargues-Fontaine curve over $S$, as defined in the discussion proceeding [7, Definition 3.3.2].
@@
  • For $b\in B(G)$, we write $\mathcal{E}_{b}$ for the associated $G$-bundle on $X$, as defined in [17, Definition 1.1]). As with the $\sigma$-centralizer $J_{b}$, this depends on a choice of representative in $G(\breve{\mathbb{Q}}_{p})$ of the $\sigma$-conjugacy class attached to $b$, but the isomorphism class does not, as a result we will suppress this technicality.
@@
  • To a diamond or $v$-stack $X$ over $\ast$, we write $\mathrm{D}(X,\overline{\mathbb{F}}_{\ell})$ for the category of étale $\overline{\mathbb{F}}_{\ell}$-sheaves, as defined in [56, Definition 14.13]. We let $\mathrm{D}^{\mathrm{ULA}}(X,\overline{\mathbb{F}}_{\ell})$ denote the full subcategory of ULA sheaves over $\ast$.
  • For an Artin $v$-stack $X$ and $\Lambda\in\{\overline{\mathbb{F}}_{\ell},\overline{\mathbb{Z}}_{\ell},\overline{\mathbb{Q}}_{\ell}\}$, we write $\mathrm{D}_{\blacksquare}(X,\Lambda)$ for the condensed $\infty$-category of solid $\overline{\mathbb{F}}_{\ell}$-sheaves on $X$, and write $\mathrm{D}_{\mathrm{lis}}(X,\Lambda)\subset\mathrm{D}_{\blacksquare}(X,\Lambda)$ for the full sub-category of $\Lambda$-lisse-étale sheaves, as defined in [16, Chapter VII].
  • If $X$ is an Artin $v$-stack ([16, Definition IV.V.1]) admitting a separated cohomologically smooth surjection $U\rightarrow X$ from a locally spatial diamond $U$ such that the étale site has a basis with bounded $\ell$-cohomological dimension (which will always be the case for our applications) then we will regard $\mathrm{D}(X,\overline{\mathbb{F}}_{\ell})$ as a condensed $\infty$-category via the identification $\mathrm{D}_{\mathrm{lis}}(X,\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(X,\overline{\mathbb{F}}_{\ell})$ when viewed as objects in $\mathrm{D}_{\blacksquare}(X,\overline{\mathbb{F}}_{\ell})$ [16, Proposition VII.6.6].
@@
  • If $F$ denotes the splitting field of $G$ then the action of $W_{\mathbb{Q}_{p}}$on $\hat{G}$ factors through $Q:=W_{\mathbb{Q}_{p}}/W_{F}$. We let $\phantom{}{}^{L}G:=\hat{G}\rtimes Q$ denote the $L$-group.
@@
[…]
  • To any condensed $\infty$-category $\mathcal{C}$, we write $\mathcal{C}^{BW^{I}_{\mathbb{Q}_{p}}}$ for the category of objects with continuous $W_{\mathbb{Q}_{p}}^{I}$-action, as defined in [16, Section IX.1].
  • For any separated $v$-stack, $X\rightarrow\mathop{\rm Spa}(K,\mathcal{O}_{K})$ where $\mathop{\rm Spa}(K,\mathcal{O}_{K})$ is a non-archimedean field, we write $\overline{X}$ for the canonical compactification of $X$ with respect to the structure map ([56, Proposition 18.6], [32, Theorem 5.15]).
  • For a reductive group $H/\mathbb{Q}_{p}$, we write $\mathrm{D}(H(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ for the (left-completed) unbounded derived category of smooth $\overline{\mathbb{F}}_{\ell}$-representations.